# Loading-dependent microscale measures control bulk properties in granular material: an experimental test of the Stress-Force-Fabric relation

### v1

@@ Abstract
The bulk behaviour of granular materials is tied to its mesoscale and particle-scale features: strength properties arise from the buildup of various anisotropic structures at the particle-scale induced by grain connectivity (fabric), force transmission, and frictional mobilization. More fundamentally, these anisotropic structures work collectively to define features like the bulk friction coefficient and the stress tensor at the macroscale and can be explained by the Stress-Force-Fabric (SFF) relationship stemming from the microscale. Although the SFF relation has been extensively verified by discrete numerical simulations, a laboratory realization has remained elusive due to the challenge of measuring both normal and frictional contact forces. In this study, we analyze experiments performed on a photoelastic granular system under four different loading conditions: uniaxial compression, isotropic compression, pure shear, and annular shear. During these experiments, we record particle locations, contacts, and normal and frictional forces vectors to measure the particle-scale response to progressing strain. We track microscale measures like the packing fraction, average coordination number and average normal force along with anisotropic distributions of contacts and forces. We match the particle-scale anisotropy to the bulk using the SFF relation, which is founded on two key principles, a Stress Rule to describe the stress tensor and a Sum Rule to describe the bulk friction coefficient; we find that the Sum and Stress Rules accurately describe bulk measurements. Additionally, we test the assumption that fabric and forces transmit load equally through our granular packings and show that this assumption is sufficient at large strain values, and can be applied to areas like rock mechanics, soft colloids, or cellular tissue where force information is inaccessible.
@@ I Introduction
Anisotropic force chain structures appear when a granular material is under load, on average aligning with the principal stress direction [7]. The presence of these mesoscale features have been used to address questions like the origin of jamming [8], plastic deformations [9] and the heterogeneous distribution of stress [10]. Recent studies have used force chains to study local stress distributions [11] and emergent elasticity [12, 13]. Not only does the loading affect the bulk properties, but it also has a lasting effect on the dynamics of the material. Granular materials in particular have the potential to store memory of loading through history dependence due to the absence of internal driving to restore the system to equilibrium [14]. Thus, the bulk properties of a granular packing under load depends on the loading geometry and the emergence of anisotropic structures (force chains and contacts) during loading to ensure the mechanical strength of the assembly.
@@
However, because the SFF relation requires particle-scale information, it has yet to be fully applied to laboratory experiments. With improvements to imaging and computing power, we are now able to use photoelastic particles to test the SFF relation in laboratory conditions and find that it is an accurate description for all of the conditions that we have applied it to. This provides a common framework with which to describe a diverse set of experimental results, and make thereby make cross-comparisons.
@@ II Methods
To test the stress-force-fabric relation at larger strains, including reaching steady state, we performed a fourth type of loading in an annular geometry as shown Fig. 1d). This quasi-two-dimensional annular shear cell, first described in [30] contains $N_{p}$ = 1810 bidisperse (also Vishay PhotoStress PSM-4, bulk modulus $E$= 4 MPa, density 1.06 g/cm3, $\mu_{p}~{}\approx 0.3$) circular disks with an equal fraction of radii $R_{s}$ = 4.5 mm and $R_{l}$ = 5.5 mm ($R_{l}/R_{s}=1.2$). The particles are dusted with baking powder as a dry lubricant to reduce basal friction between the particles and the supporting sheet. The inner and outer boundaries have particle-sized rough features to reduce slip between the boundaries and the grains. The inner boundary of the annulus is rotated at a constant rate of 1 rotation every 2.5 hours and the packing undergoes shear analogous to simple shear in an annular geometry. Using the annular strain rate $\dot{\gamma}=\frac{1}{2}\left(\frac{\partial v}{\partial R}-\frac{v}{R}\right)$ as described by Tang et al. [31], and integrating over time with appropriate boundary conditions, we calculate the strain $\epsilon$. Further details of the system and this new dataset can be found in the companion article Lee et al. [32], where we additionally investigate the long-term, transient dynamics.
@@ III.1 Stress-Force-Fabric relation
where $\vec{\ell}$ is the branch vector which joins the center of two contacting particles, $(\theta)$ indicates the average of the value within the brackets as evaluated for all contacts that are within an angular element $\Delta\theta$ centered at angle $\theta$. We note a subtlety in notation, where the summation runs over all particles in Eq.1 while it runs over all contacts in Eq.3. The average outer product between the branch vector and the force vector is multiplied by the number of contacts $\Delta M(\theta)$ that fall within the angular element when it is centered about $\theta$.
@@
$\langle f_{n}\rangle(\theta)=f_{0}\left(1+a_{n}\cos 2(\theta-\theta_{f})\right),$ (9)
@@ IV.1 Measuring angular distributions
Figure 2 shows the evolution of the different angular distributions $E(\theta)$ (top row), $\langle f_{n}\rangle(\theta)$ (middle row) and $\langle f_{t}\rangle(\theta)$ (bottom row) for each of the loading geometries as the packing is under strain (shown via the colorbar). For videos of the co-evolution of the positions, force networks, and angular histograms, see the SI videos available for each of four loading geometries. For the three rectangular geometries, $\theta$ is taken relative to the positive $\hat{x}$ coordinate, while for the annular geometry it is taken relative to the radial coordinate $R$.
Figure 2: Top row: angular distributions of the contact normals $E(\theta)$; Middle row: mean normalized normal forces $\langle f_{n}\rangle(\theta)$; Bottom row: mean normalized tangential forces $\langle f_{n}\rangle(\theta)$. Each distribution is plotted for the four loading histories (a) uniaxial, (b) isotropic, (c) pure shear and (d) annular shear. For the three rectangular geometries (a-c), $\theta$ is measured relative to the system-wide $+x$ axis, while for the annular geometry it is measured relative to the radial axis $R$. The increments of volumetric or deviatoric strain for each loading history are (a) $\epsilon=[0,0.013]$, (b) $\epsilon=[0,0.013]$ (c) $\epsilon=[0,0.06]$ and (d) $\epsilon=[0,0.3]$ in 10 evenly spaced intervals, with the value indicated by color ranging from low (dark blue) to high (light yellow).
@@
More specifically, in the case of uniaxial compression (a) and isotropic compression (b), the peaks observed on $E(\theta)$ attenuate until the distribution becomes almost isotropic $E(\theta)\sim 1$ in both cases (i.e. no preferred contact orientation is observed in the densest state). The initially isotropic shape of the normal force distribution $\langle f_{n}\rangle(\theta)$ becomes anisotropic and oriented along the loading direction in case (a) and, as expected, remains globally isotropic in case (b). In cases (a) and (b) the systems are compressed, therefore, the deformation stops when the contact network and the force network reach equilibrium.
@@
Regardless of all the subtle variations in the shape of the distributions we have described, the main result is that our experiments show a regular behaviour with $\pi$ periodicity of the angular distributions of the contacts and forces distributions. Thus, as shown in Fig.2, our experimental data for each distribution are well described by Eq.8, Eq.9 and Eq.10 at increasing levels of strain.
@@ IV.2 Texture and anisotropies
The top row of Fig. 3 shows the evolution of both coordination number $z$ and packing density $\nu$ for all four loading geometries (a–d), averaged over all trials, as a function of strain $\epsilon$. We distinguish the evolution of $z$ computed without (blue) or with (red) rattlers (i.e. particles with less than 2 contacts) and use $z$ without rattlers in Eq. 11. Initially, the values of $z$ are quite low ($z\approx 2$ measured without rattlers and $z\approx 1$ with rattlers) for all geometries tested. We observe $\nu\approx 0.73$ to $0.75$ in the rectangular geometries (a–c), consistent with our prior results in this apparatus [45], but as low as $\nu=0.64$ in the sheared annulus (d). This low value arises from significant exclusion effects near the two rough walls and heterogeneity in the packing while under shear. Under loading, $\nu$ remains globally constant (trivially in column (c), as the pure shear test is performed to conserve area). This behavior was as also reported in numerical discrete element studies of sheared frictionless (or weakly frictional, as in our experiments) grain packing [41, 42].
@@
The middle row of Fig. 3 reports the evolution of the mean normal contact force $f_{0}$ for all four loading geometries (a–d), averaged over all trials, as a function of strain $\epsilon$. We observe that $f_{0}$ increases with $\epsilon$, even as $z$ remains constant in the rectangular loading geometries (a–c). This means that the contacts support increasingly strong forces under the effects of increased strain. In contrast, for the angular shear geometry (d), the value of $f_{0}$ saturates in accordance with $z$, corresponding to the fact that the flow reaches a steady state.
Finally, the bottom row of Fig. 3 shows the evolution of the three anisotropic parameters ($a$, $a_{n}$ and $a_{t}$) for all four loading geometries, averaged over all trials, as a function of the strain $\epsilon$. First, we note that in the initial state ($\epsilon=0$), $a$ is near $0.4$ in the three of the geometries (a,b,d) as high as to $0.75$ in (c). This means that our particle-placement protocols already induce a preferential orientation of the contacts. Upon loading, in all cases, $a$ first decreases with $\epsilon$. When considered in concert with the observed increase in $z$, this highlights that more contacts are created perpendicular to the main shear direction than parallel to it. Once $z$ becomes constant, later in the loading process, different behaviours are obtained depending on the geometry tested. For geometries (a) and (b), $a$ continues to decrease, and reaches a value close to 00 for both in the large-strain limit. This reflects the fact that in this state the particles tend to reorganise without forming new contacts. In contrast, for geometries (c) and (d), $a$ tends to a plateau with an approximate value of $0.25$ for case (c) and $0.19$ for case (d), implying that the initial anisotropic nature of the contact network is preserved, albeit less pronounced.
@@ IV.3 Evaluating the bulk response via the SFF
Next, we consider the macroscopic behaviour of each loading geometry, through the evolution of the components of the stress tensor under strain. We test the Stress Rule (Eq. 11) and thereby propose a more detailed view of the role of the fabric and force network on bulk properties. Fig. 4 shows the four stress components calculated according to Eq. 1 and averaged over all trials for each geometry (a)-(d).
@@ V Conclusions
In this study, we have shown experimental validation of the SFF relation using photoelastic particles in four different loading geometries. We track the development of the coordination number, average force, and packing fraction as a function of strain along with anisotropy in the distribution of fabric and forces. The success of the SFF relation indicates that the bulk friction coefficient of a granular material arises due to a combination of anisotropy in both fabric and forces. As well, we compare the role of the anisotropies with other common characteristics and show that considering the contribution of the anistropies is necessary to fully capture the evolution of the stress tensor with increasing strain. Importantly, the SFF relation remains valid even under different geometries and loading geometry beyond the system for which the framework was initially developed. We show that there are similarities how the anisotropy amplitudes of the two compressive loadings develop under strain, something to be expected, as the uniaxial case is an extension of the isotropic compression. In addition, we have shown that the first order Fourier expansion of the stress tensor is sufficient to described bulk properties, like the bulk friction coefficient and the stress tensor components.

### v2

@@ Abstract
The bulk behaviour of granular materials is tied to its mesoscale and particle-scale features: strength properties arise from the buildup of various anisotropic structures at the particle-scale induced by grain connectivity (fabric), force transmission, and frictional mobilization. More fundamentally, these anisotropic structures work collectively to define features like the bulk friction coefficient and the stress tensor at the macroscale and can be explained by the Stress-Force-Fabric (SFF) relationship stemming from the microscale. Although the SFF relation has been extensively verified by discrete numerical simulations, a laboratory realization has remained elusive due to the challenge of measuring both normal and frictional contact forces. In this study, we analyze experiments performed on a photoelastic granular system under four different loading conditions: uniaxial compression, isotropic compression, pure shear, and annular shear. During these experiments, we record particle locations, contacts, and normal and frictional forces to measure the particle-scale response to progressing strain. We track microscale measures like the packing fraction, average coordination number and average normal force along with anisotropic distributions of contacts and forces. We match the particle-scale anisotropy to the bulk using the SFF relation, which is founded on two key principles, a Stress Rule to describe the stress tensor and a Sum Rule to describe the bulk friction coefficient; we find that the Sum and Stress Rules accurately describe bulk measurements. Additionally, we test the assumption that fabric and forces transmit load equally through our granular packings and show that this assumption is sufficient at large strain values, and can be applied to areas like rock mechanics, soft colloids, or cellular tissue where force information is inaccessible.
@@ I Introduction
Anisotropic force chain structures appear when a granular material is under load, on average aligning with the principal stress direction [7]. The presence of these mesoscale features have been used to address questions like the origin of jamming [8], plastic deformations [9] and the heterogeneous distribution of stress [10]. Recent studies have used force chains to study local stress distributions [11] and emergent elasticity [12, 13]. Not only does the loading affect the bulk properties, but it also has a lasting effect on the dynamics of the material. Granular materials in particular store memory of loading through history dependence due to the absence of internal driving to restore the system to equilibrium [14]. Thus, the bulk properties of a granular packing under load depends on the loading geometry and the emergence of anisotropic structures (force chains and contacts) during loading to ensure the mechanical strength of the assembly.
@@
However, because the SFF relation requires particle-scale information, it has yet to be fully applied to laboratory experiments. With improvements to imaging and computing power, we are now able to use photoelastic particles to test the SFF relation in laboratory conditions and find that it is an accurate description for all of the conditions that we have applied it to. This provides a common framework with which to describe a diverse set of experimental results, and thereby make cross-comparisons.
@@ II Methods
To test the stress-force-fabric relation at larger strains, including reaching steady state, we performed a fourth type of loading in an annular geometry as shown Fig. 1(d). This quasi-two-dimensional annular shear cell, first described in [30] contains $N_{p}$ = 1810 bidisperse (also Vishay PhotoStress PSM-4, bulk modulus $E$= 4 MPa, density 1.06 g/cm3, $\mu_{p}~{}\approx 0.3$) circular disks with an equal fraction of radii $R_{s}$ = 4.5 mm and $R_{l}$ = 5.5 mm ($R_{l}/R_{s}=1.2$). The particles are dusted with baking powder as a dry lubricant to reduce basal friction between the particles and the supporting sheet. The inner and outer boundaries have particle-sized rough features to reduce slip between the boundaries and the grains. The inner boundary of the annulus is rotated at a constant rate of 1 rotation every 2.5 hours and the packing undergoes shear analogous to simple shear in an annular geometry. Using the annular strain rate $\dot{\gamma}=\frac{1}{2}\left(\frac{\partial v}{\partial R}-\frac{v}{R}\right)$ as described by Tang et al. [31], and integrating over time with appropriate boundary conditions, we calculate the strain $\epsilon$. Further details of the system and this new dataset can be found in the companion article Lee et al. [32], where we additionally investigate the long-term, transient dynamics.
@@ III.1 Stress-Force-Fabric relation
where $\vec{\ell}$ is the branch vector which joins the center of two contacting particles, $(\theta)$ indicates the average of the value within the brackets as evaluated for all contacts that are within an angular element $\Delta\theta$ centered at angle $\theta$. We note a subtlety in notation, where in Eq. 1 the summation runs over all particles while it runs over all contacts in Eq. 3. The average outer product between the branch vector and the force vector is multiplied by the number of contacts $\Delta M(\theta)$ that fall within the angular element when it is centered about $\theta$.
@@
$\langle f_{n}\rangle(\theta)=f_{0}\left(1+a_{n}\cos 2\left(\theta-\theta_{f}\right)\right),$ (9)
@@ IV.1 Measuring angular distributions
Figure 2 shows the evolution of the different angular distributions $E^{c}(\theta)$ (top row), $\langle f_{n}\rangle(\theta)$ (middle row) and $\langle f_{t}\rangle(\theta)$ (bottom row) for each of the loading geometries as the packing is under strain (shown via the colorbar). For videos of the co-evolution of the positions, force networks, and angular histograms, see the SI videos available for each of four loading geometries. For the three rectangular geometries, $\theta$ is taken relative to the positive $x$ coordinate, while for the annular geometry it is taken relative to the radial coordinate $R$.
Figure 2: Top row: angular distributions of the contact normals $E^{c}(\theta)$; Middle row: mean normalized normal forces $\langle f_{n}\rangle(\theta)$; Bottom row: mean normalized tangential forces $\langle f_{n}\rangle(\theta)$. Each distribution is plotted for the four loading histories (a) uniaxial, (b) isotropic, (c) pure shear and (d) annular shear. For the three rectangular geometries (a-c), $\theta$ is measured relative to the system-wide $+x$ axis, while for the annular geometry it is measured relative to the radial axis $R$. The increments of volumetric or deviatoric strain for each loading history are (a) $\epsilon=[0,0.013]$, (b) $\epsilon=[0,0.013]$ (c) $\epsilon=[0,0.06]$ and (d) $\epsilon=[0,0.3]$ in 10 evenly spaced intervals, with the value indicated by color ranging from low (dark blue) to high (light yellow).
@@
More specifically, in the case of uniaxial compression (a) and isotropic compression (b), the peaks observed on $E^{c}(\theta)$ attenuate until the distribution becomes almost isotropic $E^{c}(\theta)\sim 1$ in both cases (i.e. no preferred contact orientation is observed in the densest state). The initially isotropic shape of the normal force distribution $\langle f_{n}\rangle(\theta)$ becomes anisotropic and oriented along the loading direction in case (a) and, as expected, remains globally isotropic in case (b). In cases (a) and (b) the systems are compressed, therefore, the deformation stops when the contact network and the force network reach equilibrium.
@@
Regardless of all the subtle variations in the shape of the distributions we have described, the main result is that our experiments show a regular behaviour with $\pi$ periodicity of the angular distributions of the contacts and forces distributions. Thus, as shown in Fig 2, our experimental data for each distribution are well described by Eq. 8, Eq. 9 and Eq. 10 at increasing levels of strain.
@@ IV.2 Texture and anisotropies
The top row of Fig. 3 shows the evolution of both coordination number $z$ and packing density $\nu$ for all four loading geometries (a)–(d), averaged over all trials, as a function of strain $\epsilon$. We distinguish the evolution of $z$ computed without (blue) or with (red) rattlers (i.e. particles with less than 2 contacts) and use $z$ without rattlers in Eq. 11. Initially, the values of $z$ are quite low ($z\approx 2$ measured without rattlers and $z\approx 1$ with rattlers) for all geometries tested. We observe $\nu\approx 0.73$ to $0.75$ in the rectangular geometries (a)–(c), consistent with our prior results in this apparatus [45], but as low as $\nu=0.64$ in the sheared annulus (d). This low value arises from significant exclusion effects near the two rough walls and heterogeneity in the packing while under shear. Under loading, $\nu$ remains globally constant (trivially in column (c), as the pure shear test is performed to conserve area). This behavior was as also reported in numerical discrete element studies of sheared frictionless (or weakly frictional, as in our experiments) grain packing [41, 42].
@@
The middle row of Fig. 3 reports the evolution of the mean normal contact force $f_{0}$ for all four loading geometries (a)–(d), averaged over all trials, as a function of strain $\epsilon$. We observe that $f_{0}$ increases with $\epsilon$, even as $z$ remains constant in the rectangular loading geometries (a)–(c). This means that the contacts support increasingly strong forces under the effects of increased strain. In contrast, for the angular shear geometry (d), the value of $f_{0}$ saturates in accordance with $z$, corresponding to the fact that the flow reaches a steady state.
Finally, the bottom row of Fig. 3 shows the evolution of the three anisotropic parameters ($a$, $a_{n}$ and $a_{t}$) for all four loading geometries, averaged over all trials, as a function of the strain $\epsilon$. First, we note that in the initial state ($\epsilon=0$), $a$ is near $0.4$ in the three of the geometries (a),(b), and (d) as high as to $0.75$ in (c). This means that our particle-placement protocols already induce a preferential orientation of the contacts. Upon loading, in all cases, $a$ first decreases with $\epsilon$. When considered in concert with the observed increase in $z$, this highlights that more contacts are created perpendicular to the main shear direction than parallel to it. Once $z$ becomes constant, later in the loading process, different behaviours are obtained depending on the geometry tested. For geometries (a) and (b), $a$ continues to decrease, and reaches a value close to 00 for both in the large-strain limit. This reflects the fact that in this state the particles tend to reorganise without forming new contacts. In contrast, for geometries (c) and (d), $a$ tends to a plateau with an approximate value of $0.25$ for case (c) and $0.19$ for case (d), implying that the initial anisotropic nature of the contact network is preserved, albeit less pronounced.
@@ IV.3 Evaluating the bulk response via the SFF
Next, we consider the macroscopic behaviour of each loading geometry, through the evolution of the components of the stress tensor under strain. We test the Stress Rule (Eq. 11) and thereby propose a more detailed view of the role of the fabric and force network on bulk properties. Fig. 4 shows the four stress components calculated according to Eq. 1 and averaged over all trials for each geometry (a)–(d).
@@ V Conclusions
In this study, we have shown experimental validation of the SFF relation using photoelastic particles in four different loading geometries. We track the development of the coordination number, average force, and packing fraction as a function of strain along with anisotropy in the distribution of fabric and forces. The success of the SFF relation indicates that the bulk friction coefficient of a granular material arises due to a combination of anisotropy in both fabric and forces. As well, we compare the role of the anisotropies with other common characteristics and show that considering the contribution of the anistropies is necessary to fully capture the evolution of the stress tensor with increasing strain. Although not considered here, we presume the shear modulus should follow a similar dependence on the evolution of the coordination number [51]. Importantly, the SFF relation remains valid even under different geometries and loading geometry beyond the system for which the framework was initially developed. We show that there are similarities how the anisotropy amplitudes of the two compressive loadings develop under strain, something to be expected, as the uniaxial case is an extension of the isotropic compression. In addition, we have shown that the first order Fourier expansion of the stress tensor is sufficient to described bulk properties, like the bulk friction coefficient and the stress tensor components.